OpenStudy (anonymous):

lim x-2+=((x+3)((x+2)/x(x+2))

OpenStudy (freckles):

cancel the common factor there that you ahve on top and bottom

OpenStudy (anonymous):

the first x+2 is an absolute value will this make a difference? I didnt see the equation options before. should read lim x 2(+)=\[\left( x+3 \right)*(\left| x+2 \right|/x+2)\]

OpenStudy (anonymous):

answer should equal 1

OpenStudy (freckles):

does make a difference

OpenStudy (anonymous):

as i suspected, didn't work canceling

OpenStudy (freckles):

\[\lim_{x \rightarrow 2^+}\frac{(x+3)|x+2|}{x(x+2)}\] is this right?

OpenStudy (freckles):

just plug in 2 because the bottom isn't going to be zero

OpenStudy (anonymous):

no the x+3 is before the equation of the \[\frac{ \left| x+2 \right|}{ x+2}\]

OpenStudy (freckles):

or did you mean -2^+

OpenStudy (anonymous):

yes -2

OpenStudy (freckles):

\[\lim_{x \rightarrow -2^+}(x+3)\frac{|x+2|}{x+2}\] so there is no x on bottom I guess? anyways we have x approaches -2 from the right that means x is greater than -2 x>-2 adding 2 on both sides gives x+2>0 so since x+2>0 then |x+2|=x+2 so you can actually cancel the x+2's out (since we are approaching 2 from right)

OpenStudy (freckles):

-2 from right*

OpenStudy (anonymous):

that makes sense but it still doesn't get the answer 1 thouhh

OpenStudy (freckles):

yes

OpenStudy (freckles):

it does

OpenStudy (freckles):

-2+3 is indeed 1

OpenStudy (anonymous):

oh yes thank you, sometimes the easiest stuff hangs me up.

OpenStudy (freckles):

\[\lim_{x \rightarrow -2^-}(x+3)\frac{|x+2|}{x+2} \] if we had approaches -2 from left then we know x<-2 adding 2 on both sides would give us x+2<0 so since we know x+2<0 then |x+2|=-(x+2)

OpenStudy (freckles):

so here you would have (-2+3)(-1)

OpenStudy (freckles):

anyways since right limit=1 and left limit=-1 then the actual limit wouldn't exist

OpenStudy (freckles):

just extra work if you wanted more examples

OpenStudy (anonymous):

i appreciate the help, that does make sense there is no actual limit for the whole equation even though you can figure the left and right sides.

OpenStudy (freckles):

right :) if the left=right then the actual would equal whatever that was they equaled :)